twice the sum of a number and 3 is equal to three times the difference of the number and 2. Find the number
Answer by checkley77(12569)

2(N+3)=3(N-2)
2N+6=3N-6
2N-3N=-6-6
-N=-12
N=12 ANSWER.
PROOF:
2(12+3)=3(12-2)
2*15=3*10
30=30